A sign (mass 2300 kg ) hangs from the end of a vertical steel girder with a cross-sectional area of 0.018 m2 . Ignore the mass of the girder itself. The elastic modulus for steel is 2.0×10^11N/m2.

A.) What is the stress within the girder?

B.)What is the strain on the girder?
C.)If the girder is 9.40 m long, how much is it lengthened?
Answer #1

A] stress = F/A = mg/A = 2300*9.8/0.018 = 1252222 N/m^2

B] Strain = stress/Y = 1252222/2e11 = 0.00000626

C] delta L = L*strain = 9.4*0.00000626 = 0.0000588 m